OpenStudy (anonymous):

factor completely 2x^4-72

OpenStudy (anonymous):

Set the factor '(-6 + x2)' equal to zero and attempt to solve: Simplifying -6 + x2 = 0 Solving -6 + x2 = 0 Move all terms containing x to the left, all other terms to the right. Add '6' to each side of the equation. -6 + 6 + x2 = 0 + 6 Combine like terms: -6 + 6 = 0 0 + x2 = 0 + 6 x2 = 0 + 6 Combine like terms: 0 + 6 = 6 x2 = 6 Simplifying x2 = 6 Take the square root of each side: x = {-2.449489743, 2.449489743} Solution x = {-2.449489743, 2.449489743}

OpenStudy (anonymous):

^^^^^^^ this is after the process of simplifing